# Colimits of Heyting Algebras through Esakia Duality

### v1

@@ 3 Free Constructions of Heyting Algebras over Distributive Lattices
In this section we explain how one can construct a Heyting algebra freely from any given distributive lattice. This construction generalises the case of Heyting algebras freely generated by finite distributive lattices, as originally analyzed by Ghilardi [16]. The key tools at hand will be the Priestley/Esakia duality of distributive lattices.
3.1 Conceptual Idea of the Construction
Before diving into the details, we will provide an informal explanation of the construction by Ghilardi which justifies the technical developments, following the core of the discussion from [7]. We urge the reader to consult this section for intuition whilst going through the details of the next section.
Suppose that $D$ is a finite distributive lattice, and that we wish to construct a Heyting algebra from $D$. Then first we must (1) freely add to $D$ implications of the form $a\rightarrow b$ for $a,b\in D$. This is the same as freely generating a distributive lattice out of $D^{2}$, and considering its coproduct with $D$. Dually, if $X$ is the dual poset to $D$, this will then be the same as considering
@@ 3.1 Conceptual Idea of the Construction
given that the free distributive lattice on a set of pairs of generators $X$ will be dual to $\mathcal{P}(X\times X)$ (see Section 2). However, we need (2) to impose axioms forcing these implications to act like relative pseudocomplements. A first move is to impose the axiom of a Weak Heyting Algebra:
@@
Applying a quotient under these axioms, will dually yield (see [7, Theorem 3.5]),
@@
The fact that we want to obtain a genuine relative pseudocomplement, means we need to impose further axioms, which implies throwing out some of the elements from $\mathcal{P}(X)$. The additional fact that we would want the map $i_{0}$ to be a distributive lattice homomorphism suggests a way of doing this: make it so that for each pair of upsets $U,V$, $[U]\cup[V]=[U\cup V]$. After some thought one can then be lead to consider rooted subsets of $X$, obtaining the set $\mathcal{P}_{r}(X)$. Some verifications show that this will indeed be the poset that is needed: that
@@
Now at this point we will have added all implications to elements $a,b\in D$, obtaining a distributive lattice $D_{1}$, but all the new implications added might not in turn have implications between themselves. So we need to (3) iterate the construction, infinitely often, to add all necessary implications. However, the final complication is that each step of this construction adds implications to every element in the previous lattice. So in particular, $D_{2}$ will contain a fresh relative pseudocomplement, $[a]\rightarrow[b]$. If we let the construction run infinitely often in this way, it could be that in the end no element would be the relative pseudocomplement of $a$ and $b$, so we need to ensure that on the second iteration, the previously added relative pseudocomplements are preserved, i.e.:
@@
as will be explained in the next section, where $g$ is some order-preserving map which serves to index the relative pseudocomplements which are to be preserved.
With this in mind, we turn in the next section to the technical details involved in this generalization.
@@ 3.2 G-Open Subsets and Vietoris Functors
We begin by adapting the notion of $g$-openness to our setting. Let $X,Y$ be two Priestley spaces. We recall that a Priestley morphism $f:X\to Y$ is an order-preserving and continuous map. By Priestley duality, such maps are dual to distributive lattice homomorphisms $f^{-1}:\mathsf{ClopUp}(Y)\to\mathsf{ClopUp}(X)$.
@@
Let $X,Y,Z$ be Priestley spaces, and $f:X\to Y$ and $g:Y\to Z$ be Priestley morphisms. We say that $f$ is open relative to $g$ ($g$-open for short) if $f^{-1}$ preserves relative pseudocomplements of the kind $g^{-1}[U]\rightarrow g^{-1}[V]$ where $U,V\in\mathsf{ClopUp}(Z)$.
@@
Asssume that $f$ satisfies (*). We want to show that $f^{-1}[g^{-1}[U]\rightarrow g^{-1}[V]]=f^{-1}g^{-1}[U]\rightarrow f^{-1}g^{-1}[V]$. Note that the left to right inclusion always holds because $f$ is a Priestley morphism. So assume that $x\in f^{-1}g^{-1}[U]\rightarrow f^{-1}g^{-1}[V]$. Suppose that $f(x)\leq y$, and $y\in g^{-1}[U]$. By assumption, there is some $x^{\prime}$ such that $x\leq x^{\prime}$ and $g(f(x^{\prime}))=g(y)$; hence $f(x^{\prime})\in g^{-1}[U]$, so by assumption, $x^{\prime}\in f^{-1}g^{-1}[U]$, and so, $x^{\prime}\in f^{-1}g^{-1}[V]$. This means that $f(x^{\prime})\in g^{-1}[V]$, so $y\in g^{-1}[V]$. This shows what we want.
@@
Now given $g:X\to Y$ a Priestley morphism, and $S\subseteq X$ a closed subset, we say that $S$ is g-open (understood as a poset with the restricted partial order relation) if the inclusion is itself $g$-open. This means by the above lemma that $S$ is $g$-open if the following condition holds:
@@
At this stage we would like to point out to the reader that the fact that we pick a given unique $g$ is totally incidental; in fact, whilst the results of this section will be proved for a single $g$, one could take any number of continuous morphisms and obtain the same results. We will not pursue such considerations further in this paper, however.
Throughout, fix $g:X\to Y$ a continuous and order-preserving map (the case for several such maps being preserved is entirely similar). Recall from Subsection 2.2 that $(V(X),\preceq)$ is a Priestley space. From this space we can move closer to the space we will be interested, by first considering $V_{r}(X)\subseteq V(X)$, the space of rooted closed subsets, with the induced order and the subspace topology. On this we can prove the following [^2] This is stated without proof in [8, Lemma 6.1]. The key idea of the proof below was communicated to me by Mamuka Jibladze, to whom I am greatly indebted.:
@@
We note that $\ni_{X}$ is closed, since it is the pullback of the continuous map {}$\{\}:X\to V(X)$ along the second projection of $\leq\subseteq V(X)\times V(X)$ (which is a closed relation, since $V(X)$ is Priestley by Lemma 6). On the other hand, $LB(X)$ can also be seen to be closed, through a direct argument: suppose that $(C,r)\notin LB(X)$. By definition, then, $C\nsubseteq{\uparrow}r$, so there is a point $y\in C$ such that $r\nleq y$. By the Priestley separation axiom, there is a clopen upset $U$ such that $r\in U$ and $y\notin U$. So consider
@@
Finally, we will now refine the rooted subsets to the ones we are truly interested in – the $g$-open ones. The following lemma encapsulates the key fact about this.
@@
Assume that $M$ is not $g$-open. This means that if $x$ is the root of $M$, there is some $y$ such that $x\leq y$, but for each $k\in M$, $g(k)\neq g(y)$. Hence $g(y)\notin g[M]$, and hence $\{g(y)\}\cap g[M]=\emptyset$. Since $g$ is a continuous map between Stone spaces, $g[M]$ is closed; hence, since this is a Priestley space, there is a clopen set of the form $-U_{i}\cup V_{i}$ for $U_{i},V_{i}$ clopen upsets, such that $g[M]\subseteq-U_{i}\cup V_{i}$, and $g(y)\notin-U_{i}\cup V_{i}$. Then we have that $M\subseteq g^{-1}[-U_{i}]\cup g^{-1}[V_{i}]$ and $M\cap{\downarrow}(g^{-1}[U_{i}]-g^{-1}[V_{i}])\neq\emptyset$. Hence consider
$S=[g^{-1}[-U_{i}]\cup g^{-1}[V_{i}]]\cap\langle{\downarrow}(g^{-1}[U_{i}]-g^{-1}[V_{i}])\rangle.$
Because $X$ has relative pseudocomplements of sets of the form $g^{-1}[U]$, by Proposition 4 we have that ${\downarrow}(g^{-1}[U_{i}]-g^{-1}[V_{i}])$ will be clopen in $X$. Hence the set $S$ is an intersection of subbasic sets. Moreover, $M$ belongs there. Now if $N$ belongs there, then first $N\subseteq g^{-1}[-U_{i}\cup V_{i}]$; and also, by downwards closure, its root $k$ will be below some point $y^{\prime}$ in $g^{-1}[U_{i}]-g^{-1}[V_{i}]$. But then for each $m\in N$, $g(m)\neq g(k)$, so $N$ is not $g$-open.
@@
We note that the hypothesis of the previous lemma – that $X$ will have relative pseudocomplements of the form $g^{-1}[U]\rightarrow g^{-1}[V]$ – will always be satisfied in our contexts: this is how we think of $g$ a way to parametrise those pseudocomplements which we wish to preserve.
@@ 3.3 Free Heyting Algebras from Distributive Lattices
We will now put the tools developed in the previous section to use.
@@
Let $g:X\to Y$ be a Priestley morphism. The $g$-Vietoris complex over $X$ $(V_{\bullet}^{g}(X),\leq_{\bullet})$, is a sequence
@@
Note that given a $g$-Vietoris complex, one can form the projective limit of the sequence, which, since Priestley spaces are closed under such a construction, will again be a Priestley space. We denote this limit by $V_{G}^{g}(X)$ (the $G$ standing for “Ghilardi”).
The purpose of such a construction lies in the universal property which is carried out over any one step, which we proceed to outline:
@@
{tikzcd}
@@
(Of Lemma 13) We show the property from Lemma 14. Given $Z,X$ Priestley spaces as given, the definition of $h^{\prime}$ is totally forced by the commutativity of the triangle. Indeed, given $a\in Z$, let
$h^{\prime}(a)=\{h(b):a\leq b\}.$
Note that $h^{\prime}(a)=h[{\uparrow}a]$, and since $h$ is continuous between Stone spaces, and hence a closed map, $h^{\prime}$ maps $a$ to a closed subset; it is of course also rooted. And it is $g$-open as a subset, because $h$ is $g$-open as a map. Clearly $r_{g}(h^{\prime}(a))=h(a)$. It is continuous, since $(h^{\prime})^{-1}[[-U\cup V]]=\{a:h^{\prime}(a)\in[-U\cup V]\}$, and this is the same as saying that ${\uparrow}a\subseteq h^{-1}[-U]\cup h^{-1}[V]$. Now since $Z$ has relative pseudocomplements of these sets, then $\Box(-U\cup V)$ exists, and indeed
@@
This shows continuity. It is also of course order-preserving, since if $a\leq b$, then ${\uparrow}b\subseteq{\uparrow}a$, so $h[{\uparrow}b]\subseteq h[{\uparrow}a]$, i.e., $h^{\prime}(a)\leq h^{\prime}(b)$. Finally it is also $r_{g}$-open: if $a\in Z$, and $h^{\prime}(a)\leq M$ where $M\in V_{g}(X)$, then $M\subseteq h^{\prime}(a)$, so if $z$ is the root of $M$, then $z=h(c)$, where $a\leq c$. Then $r(h^{\prime}(c))=r(M)$, as desired. The uniqueness is enforced by the shape of the diagram.∎
All of this allows us to prove the following key theorem:
@@
Let $\mathcal{A}$ be a Heyting algebra, let $\mathcal{D}$ be a distributive lattice, and let $p:\mathcal{A}\to\mathcal{D}$ be a distributive lattice homomorphism which preserves the relative pseudocomplements from $\mathcal{A}$ (such that $D$ contains them). Let $\mathcal{H}$ be the dual of $V_{G}^{p}(X)$. Then $\mathcal{H}$ is the Heyting algebra freely generated by $D$ which preserves the relative pseudocomplements coming from $p$.

### v3

@@ 3 Free Constructions of Heyting Algebras over Distributive Lattices
In this section we explain how one can construct a Heyting algebra freely from any given distributive lattice. This construction generalises the case of Heyting algebras freely generated by finite distributive lattices, as originally analyzed by Ghilardi [ghilardifreeheyting].
3.1 Conceptual Idea of the Construction and the Finite Case
Before diving into the details, we will provide an informal explanation of the construction by Ghilardi which justifies the technical developments, following the core of the discussion from [Bezhanishvili2011]. We urge the reader to consult this section for intuition whilst going through the details of the next section.
Suppose that $D$ is a finite distributive lattice, and that we wish to construct a Heyting algebra from $D$. Then certainly we must (1) freely add to $D$ implications of the form $a\rightarrow b$ for $a,b\in D$. This is the same as freely generating a distributive lattice out of $D^{2}$, and considering its coproduct with $D$. Dually, if $X$ is the dual poset to $D$, this will then be the same as considering
@@ 3.1 Conceptual Idea of the Construction and the Finite Case
given that the free distributive lattice on a set of pairs of generators $X$ is dual to $\mathcal{P}(X\times X)$ (see Section 2). However, we need (2) to impose axioms forcing these implications to act like relative pseudocomplements. A first move is to impose the axiom of a Weak Heyting Algebra:
@@
Applying a quotient under these axioms, will dually yield that from our $X$, we obtain (see [Bezhanishvili2011, Theorem 3.5]),
@@
The fact that we want to obtain a genuine relative pseudocomplement, means we need to impose further axioms, which implies throwing out some of the elements from $\mathcal{P}(X)$. The additional fact that we would want the map $i_{0}$ to be a distributive lattice homomorphism suggests a way of doing this: make it so that for each pair of upsets $U,V$, $[U]\cup[V]=[U\cup V]$. After some thought one is lead to consider rooted subsets of $X$, obtaining the poset $\mathcal{P}_{r}(X)$. Some verifications show that this will indeed be the poset that is needed: that
@@
Now at this point we will have added all implications to elements $a,b\in D$, obtaining a distributive lattice $D_{1}$, but all the new implications added might not in turn have implications between themselves. So we need to (3) iterate the construction, infinitely often, to add all necessary implications. However, the final complication is that each step of this construction adds implications to every element in the previous lattice. So in particular, $D_{2}$ will contain a fresh relative pseudocomplement, $[a]\rightarrow[b]$, which need not agree with the relative pseudocomplement $[a\rightarrow b]$. If we let the construction run infinitely often in this way, it could then be that in the end no element would be the relative pseudocomplement of $a$ and $b$, so we need to ensure that on the second iteration, the previously added relative pseudocomplements are preserved, i.e.:
@@
where $g$ is some order-preserving map which serves to index the relative pseudocomplements which are to be preserved.
To conclude this section, let us take a look at an example calculation:
Example 10.
Let $X=\mathbf{2}$ the $0<1$ poset. Then we can illustrate the first four steps of the step-by-step construction of Rieger-Nishimura ladder (see Figure 1) in Figure 2. Note that the step-by-step construction does not in general agree with the construction we provided before.
[…]
Figure 2: Four steps of the construction of the Rieger-Nishimura
@@ 3.2 G-Open Subsets and Vietoris Functors
We begin by adapting the notion of $g$-openness to our setting. Let $X,Y$ be two Priestley spaces, and let $f:X\to Y$ be a monotone and continuous map. By Priestley duality, such maps are dual to distributive lattice homomorphisms $f^{-1}:\mathsf{ClopUp}(Y)\to\mathsf{ClopUp}(X)$.
@@
Let $X,Y,Z$ be Priestley spaces, and $f:X\to Y$ and $g:Y\to Z$ be monotone and continuous functions. We say that $f$ is open relative to $g$ ($g$-open for short) if $f^{-1}$ preserves relative pseudocomplements of the kind $g^{-1}[U]\rightarrow g^{-1}[V]$ where $U,V\in\mathsf{ClopUp}(Z)$.
@@
Asssume that $f$ satisfies (*). We want to show that $f^{-1}[g^{-1}[U]\rightarrow g^{-1}[V]]=f^{-1}g^{-1}[U]\rightarrow f^{-1}g^{-1}[V]$. Note that the left to right inclusion always holds because $f$ is a monotone map. So assume that $x\in f^{-1}g^{-1}[U]\rightarrow f^{-1}g^{-1}[V]$. Suppose that $f(x)\leq y$, and $y\in g^{-1}[U]$. By assumption, there is some $x^{\prime}$ such that $x\leq x^{\prime}$ and $g(f(x^{\prime}))=g(y)$; hence $f(x^{\prime})\in g^{-1}[U]$, so by assumption, $x^{\prime}\in f^{-1}g^{-1}[U]$, and so, $x^{\prime}\in f^{-1}g^{-1}[V]$. This means that $f(x^{\prime})\in g^{-1}[V]$, so $y\in g^{-1}[V]$. This shows what we want.
@@
Now given $g:X\to Y$ a monotone and continuous function, and $S\subseteq X$ a closed subset, we say that $S$ is g-open (understood as a Priestley space with the induced order and topology) if the inclusion is itself $g$-open. This means by the above lemma that $S$ is $g$-open if the following condition holds:
@@
Remark 3.1.
The fact that we pick a given unique $g$ is totally incidental; in fact, whilst the results of this section will be proved for a single $g$, one could take any number of continuous morphisms and obtain the same results. We will need this briefly in the description of coproducts made below.
Throughout, fix $g:X\to Y$ a continuous and order-preserving map (the case for several such maps being preserved is entirely similar). Recall from Lemma 9 that $(V(X),\preceq)$ is a Priestley space. From this space we can move closer to the space we will be interested, by first considering $V_{r}(X)\subseteq V(X)$, the space of rooted closed subsets, with the induced order and the subspace topology. On this we can prove the following [^2] This is stated without proof in [Bezhanishvili2011heytingcoalgebra, Lemma 6.1]. The key idea of the proof below was communicated to me by Mamuka Jibladze.:
@@
We note that $\ni_{X}$ is closed [^3] Indeed, one could simply observe that this is the pullback of the continuous map {}$\{\}:X\to V(X)$ along the second projection of $\leq\subseteq V(X)\times V(X)$.: if $(C,x)$ is such that $x\notin C$, then $\{x\}\cap C=\emptyset$. Since $X$ is a Stone space, let $U$ be a clopen subset separating them i.e., $x\in U$ and $C\subseteq X-U$. Then we can consider,
$[X-U]\times U$
which is a neighbourhood of $(C,x)$; and certainly if $(D,y)\in[X-U]\times U$, then $y\notin D$.
In turn, $LB(X)$ can also be seen to be closed, through a direct argument: suppose that $(C,r)\notin LB(X)$. By definition, then, $C\nsubseteq{\uparrow}r$, so there is a point $y\in C$ such that $r\nleq y$. By the Priestley separation axiom, there is a clopen upset $U$ such that $r\in U$ and $y\notin U$. So consider
@@
Finally, we will refine the rooted subsets to the ones we are truly interested in – the $g$-open ones. The following lemma encapsulates the key fact about this.
@@
Assume that $M$ is not $g$-open. This means that if $x$ is the root of $M$, there is some $y$ such that $x\leq y$, but for each $k\in M$, $g(k)\neq g(y)$. Hence $g(y)\notin g[M]$, and hence $\{g(y)\}\cap g[M]=\emptyset$. Since $g$ is a continuous map between Stone spaces, $g[M]$ is closed; hence, since this is a Priestley space, there is a clopen set of the form $-U_{i}\cup V_{i}$ for $U_{i},V_{i}$ clopen upsets, such that $g[M]\subseteq-U_{i}\cup V_{i}$, and $g(y)\notin-U_{i}\cup V_{i}$. Then we have that $M\subseteq-g^{-1}[U_{i}]\cup g^{-1}[V_{i}]$ and $M\cap{\downarrow}(g^{-1}[U_{i}]-g^{-1}[V_{i}])\neq\emptyset$. Hence consider
$S=[-g^{-1}[U_{i}]\cup g^{-1}[V_{i}]]\cap\langle{\downarrow}(g^{-1}[U_{i}]-g^{-1}[V_{i}])\rangle.$
Because $X$ has relative pseudocomplements of sets of the form $g^{-1}[U]$, by Proposition 5 we have that ${\downarrow}(g^{-1}[U_{i}]-g^{-1}[V_{i}])$ will be clopen in $X$. Hence the set $S$ is an intersection of subbasic sets. Moreover, $M$ belongs there. Now if $N$ belongs there, then first $N\subseteq-g^{-1}[U_{i}]\cup g^{-1}[V_{i}]$; and also, by downwards closure, its root $k$ will be below some point $y^{\prime}$ in $g^{-1}[U_{i}]-g^{-1}[V_{i}]$. But then for each $m\in N$, $g(m)\neq g(y^{\prime})$, so $N$ is not $g$-open.
@@
We note that the hypothesis of the previous lemma – that $X$ will have relative pseudocomplements of the form $g^{-1}[U]\rightarrow g^{-1}[V]$ – will always be satisfied in our contexts. This is the sense in which we will think of $g$ as a way to parametrise those pseudocomplements which we wish to preserve.
@@ 3.3 Free Heyting Algebras from Distributive Lattices
We will now put the tools developed in the previous section together:
@@
Let $g:X\to Y$ be a monotone and continuous function between Priestley spaces $X$ and $Y$. The $g$-Vietoris complex over $X$ $(V_{\bullet}^{g}(X),\leq_{\bullet})$, is an infinite sequence
@@
Given a $g$-Vietoris complex over $X$, one can form the projective limit of the sequence, which, since Priestley spaces are closed under such a construction, will again be a Priestley space. We denote this limit by $V_{G}^{g}(X)$ (the $G$ standing for “Ghilardi”). The purpose of such a construction lies in the universal property which is carried out over any one step, which we proceed to outline:
@@
${\mathsf{ClopUp}(X)}$${\mathsf{ClopUp}(V_{g}(X))}$${D}$$\scriptstyle{\mu}$$\scriptstyle{r_{g}^{-1}}$$\scriptstyle{\mu^{\prime}}$
@@
(Of Lemma 17) We show the property from Lemma 18. Given $Z,X$ Priestley spaces as given, first we note that the definition of $h^{\prime}$ is totally forced by the commutativity of the triangle: if $k:Z\to V_{g}(X)$ is any map in these conditions, we will show that
$k(a)=h[{\uparrow}a].$
Indeed, since the diagram commutes, the root of $k(a)$ will be $h(a)$; and if $a\leq b$, then $k(a)\leq k(b)$, so $h[{\uparrow}a]\subseteq k(a)$. Now if $x\in k(a)$, consider $k(a)\cap{\uparrow}x$. Then $k(a)\leq k(a)\cap{\uparrow}x$, so because $k$ is $r_{g}$-open, there is some $a^{\prime}\geq a$ such that $r_{g}(k(a^{\prime}))=r_{g}(k(a)\cap{\uparrow}x)=x$. Hence $h(a^{\prime})=x$, and so $x=h(a^{\prime})\in h[{\uparrow}a]$, as desired.
It thus suffices to show that given $a\in Z$, the assignment
$h^{\prime}(a)=\{h(b):a\leq b\},$
indeed defines a function as desired. Note that $h^{\prime}(a)=h[{\uparrow}a]$, and since $h$ is continuous between Stone spaces, and hence a closed map, $h^{\prime}$ maps $a$ to a closed subset; it is of course also rooted and $h(a)$. And it is $g$-open as a subset, because $h$ is $g$-open as a map. Clearly $r_{g}(h^{\prime}(a))=h(a)$. It is continuous, since $(h^{\prime})^{-1}[[-U\cup V]]=\{a:h^{\prime}(a)\in[-U\cup V]\}$, and this is the same as saying that ${\uparrow}a\subseteq h^{-1}[-U]\cup h^{-1}[V]$. Now since $Z$ has relative pseudocomplements of these sets, then $\Box(-U\cup V)$ exists, and indeed
@@
This shows continuity. It is also of course order-preserving, since if $a\leq b$, then ${\uparrow}b\subseteq{\uparrow}a$, so $h[{\uparrow}b]\subseteq h[{\uparrow}a]$, i.e., $h^{\prime}(a)\leq h^{\prime}(b)$. Finally it is also $r_{g}$-open: if $a\in Z$, and $h^{\prime}(a)\leq M$ where $M\in V_{g}(X)$, then $M\subseteq h^{\prime}(a)$, so if $z$ is the root of $M$, then $z=h(c)$, where $a\leq c$. Then $r(h^{\prime}(c))=r(M)$, as desired.∎
Moreover, we can see that the construction $V_{g}$ also works functorially:
Lemma 19.
Let $p:X\to Y$ be a monotone and continuous function between Priestley spaces, and $g_{X}:X\to Z$ and $g_{Y}:Y\to Z$ be two maps such that the obvious triangle commutes, and the relative pseudocomplements indexed by $g_{X}$ and $g_{Y}$ exist. Then $p[-]:V_{g}(X)\to V_{g}(Y)$ is a unique monotone and continuous function such that for each $C\in V_{g_{x}}(X)$, $pr(C)=r_{g_{Y}}p[C]$.
Proof.
It is clear that $p[-]$ will be a monotone and continuous function. To see its uniqueness, note that $p[-]$ is $r_{g_{Y}}$-open, and hence by Lemma 17, it is the unique extension of $pr$ (which is $g_{Y}$-open given that $p$ is $g_{Y}$-open). ∎
Recall that since $V_{G}(X)$ is a projective limit, there is a map $\pi_{0}:V_{G}(X)\to X$ which projects an inverse limit to $X$. Note that $\pi_{0}$ is actually surjective, since for each $x\in X$, we can consider the sequence:
$(x,{\uparrow}x,{\uparrow}({\uparrow}x),...)$
which will be an element in $V_{G}(X)$, mapping to $x$.
Lemma 20.
Let $g:X\to Y$ be a monotone and continuous function such that $g$-indexed relative pseudocomplements exist. Suppose that $Z$ is an Esakia space and $h:Z\to X$ is a monotone and continuous function which is $g$-open. Then there exists a unique p-morphism $\overline{h}:Z\to V_{G}(X)$ such that $\pi_{0}\overline{h}=h$.
Proof.
Using repeatedly Lemma 17, given that $Z$ is an Esakia space, we get a sequence of maps, starting with $h_{0}=h$, and letting $h_{n+1}=h_{n}[-]$. Then for each $x\in Z$, we have that $\overline{h}(x)$ will be the element
$(h_{0}(x),h_{1}(x),...,h_{n}(x),...).$
It is clear that this map is an order-preserving map and continuous, by the universal property of inverse limits. We check that it is a p-morphism. Assume that for $x\in Z$, $\overline{h}(x)\leq y$. Then consider the following set:
${\uparrow}x\cap\bigcap\{h_{n}^{-1}[y(n)]:n\in\omega\}.$
Since $Z$ is compact, and all the sets involved are closed, this intersection will be empty if and only if a finite intersection is empty. Let $n$ be such that
${\uparrow}x\cap h_{0}^{-1}[y(0)]\cap...\cap h_{n}^{-1}[y(n)]=\emptyset.$
Since $\overline{h}(x)\leq y$, then note that $h_{n+1}(x)\leq y(n+1)$; so there is some $k\geq x$ such that $h_{n}(k)=y(n)$. Then certainly $k\in{\uparrow}x\cap h_{0}^{-1}[y(0)]\cap...\cap h_{n}^{-1}[y(n)]$, which contradicts the above set being empty. By reductio we get $x\leq k$ such that $\overline{h}(k)=y$, as desired. ∎
Finally, we can lift monotone and continuous functions satisfying appropriate $g$-openness conditions functorially:
Lemma 21.
Let $p:X\to Y$ be a monotone and continuous function between Priestley spaces which is $g_{Y}$-open, where $g_{X}:X\to Z$ and $g_{Y}:Y\to Z$ are two maps such that the obvious triangle commutes and the relative pseudocomplements indexed by $g_{X}$ and $g_{Y}$ exist. Then there is a unique p-morphism $\overline{p}:V_{G}(X)\to V_{G}(Y)$ such that $\pi_{0}^{Y}\overline{p}=p\pi_{0}^{X}$.
Proof.
We use Lemma 19 repeatedly on the $g_{X}$-Vietoris complex over $X$ and the $g_{Y}$-Vietoris complex over $Y$: we let $p_{0}=p$, and $p_{n+1}=p[-]$. This lifts to a function on the inverse limits mapping $x\in V_{G}(X)$ to
$(p_{0}(x(0)),p_{1}(x(1)),...,p_{n}(x(n)),...).$
It is not difficult to see that having $\overline{p}$ so defined, $\pi_{0}\overline{p}$ is a $g_{Y}$-open map, so $\overline{p}$ will additionally coincide with the unique p-morphism lifting of $p\pi_{0}$ (by Lemma 20), in addition to making all diagrams commute. ∎
We can now define $V_{G}:\mathbf{Pries}\to\mathbf{Esa}$ as a functor, assigning $V_{G}(X)$ to each Priestley space $X$, and $\overline{p}:V_{G}(X)\to V_{G}(Y)$ to each monotone and continuous function $p:X\to Y$. We then obtain:
@@
The functor $V_{G}$ is right adjoint to the inclusion $I:\mathbf{Esa}\to\mathbf{Pries}$. Moreover,